please help:
If 2x=20-(8+16) then find the value of y here
-0717y+733825=x

Answer by Rhodium(14)

First find x:
2x=20-(8+16)
2x=20-24
2x=-4
x=-2
Substitute x:
-0717y+733825=-2
-0717y=-2-733825
y=(-2-733825)/-0717
y=1023.468619